OpenStudy (anonymous):

3x-2y=4 2x-5y=21 if the system cannot be solved state that it cannot be solved and explain why.. plz help!!

jhonyy9 (jhonyy9):

3x-2y=4 multiplie by -2 2x-5y=21 multiplie by 3 -6x+4y=-8 6x-15y=63 ------------- assum equations 0 - 11y =55 y=55/(-11) = -5 so from this what yign that this system has solutions sure

OpenStudy (anonymous):

omg thank you for explaining!!

jhonyy9 (jhonyy9):

was my pleasure yw good luck bye

jhonyy9 (jhonyy9):

so after this you can calcule the x ,yes ?